There are two truck kinds available to a transportation company: Type A and Type B. When comparing Type A with Type B, Type A has an overall volume of 20 m3 with equal portions for refrigerated and non-refrigerated stock, while Type B has an overall volume of 40 m3 with equal sections for both types of stock. 3,000 m3 of refrigerated stock and 4 000 m3 of nonrefrigerated goods must be transported by truck. Type A cars cost P30 per kilometer, whereas Type B cars cost P40 per kilometer. In order to keep rental costs to a low, how many of each kind of truck should the grocer rent?


1
Expert's answer
2022-03-19T02:41:47-0400

Lets truck type A - "x" , truck type B - "y" .

Type A has an overall volume of 20 m3 with equal portions for refrigerated (10 m3 ) and non-refrigerated stock (10 m3).

Type B has an overall volume of 40 m3 with equal for refrigerated (20 m3 ) and non-refrigerated stock (20 m3).

3000 m3 of refrigerated stock

"10x+20y\geq3000"

4 000 m3 of nonrefrigerated stock

"10x+20y\geq4000"

Type A cars cost P30 per kilometer, Type B cars cost P40 per kilometer.

"F(x,y)=30x+40y\to min"

"10x+20y\geq3000\\
10x+20y\geq4000\\
F(x,y)=30x+40y\to min\\
x+2y\geq300\\
x+2y\geq400\\
x\geq0,y\geq0, x\in N, y\in N\\
x+2y\geq400\\
x+2y=400\\
\begin{matrix}
 x& 0 &200\\
 y& 20&100
\end{matrix}\\
30x+40y=0\\
\begin{matrix}
 x& 0 &4\\
 y& 0&-3
\end{matrix}\\
\vec{a}=(30,40)"

In parallel, transfer the line "30x+40y=0" in the direction of the vector "\vec{a}=(30,40)" . We are looking for a point of intersection with a straight line "x+20y=400"


We have a point (109.91, 145,455), "x\in N, y\in N\implies" "x=110, y=146"